# Have We Scene It All? Scene Graph-Aware Deep Point Cloud Compression

### v1

@@ IV Experimental Evaluation
We evaluate our method’s capability to efficiently and accurately compress a point cloud scan. We compare against the commonly used codecs: Draco [8] and the octree-based compression algorithm from Mekuria et al. [9], which we refer to as “MPEG”. Additionally we include the current state-of-the-art approaches RENO [19] and OctAttention [10], as well as RecNet [21], that in contrast to ours, encodes each scan into a single latent vector through a range-image embedding. We note that the last three methods, unlike the proposed approach, or Draco/MPEG, do not encode the semantic classes.
@@ IV-A Experimental Setup
We benchmark our method on the SemanticKITTI [27] dataset. In order to create the semantic scene graph we use the ground truth semantic labels and in addition we crop each point cloud to 50 meters for all methods, following common practices in the field [28]. Similar to SegMap [5] we use sequences 05 and 06 for training and the rest (00 to 04 and 07 to 10) for the evaluation and comparison to the baselines.
Evaluation Metrics: The quality of a compression algorithm typically involves a trade-off between compression ratio and reconstruction error. To quantify compression, we use the average bits-per-point (bpp) required to store the encoded point cloud. For our method, this includes all latent vectors as well as the semantic scene graph. Following the evaluation protocol in [13], we assess reconstruction error using three metrics. The first is the symmetric point distance $D_{\text{CD}}$, defined earlier in Eq. (6) and Eq. (7), which measures the average nearest-neighbor distance between the reconstructed and ground truth point clouds. However, for many robotic tasks that rely on point-to-plane ICP registration, reconstructing the exact same points is less critical than ensuring that points lie on the correct surfaces. To address this, we also report the symmetric plane distance $D_{\perp}$, which is computed similarly to $D_{\text{CD}}$ but replaces the Euclidean distance $d(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ with the point-to-plane distance ${d}_{n}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=|\boldsymbol{n}^{\mathsf{T}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})|$, where $\boldsymbol{n}\in\mathbb{R}^{3}$ is the ground truth surface normal at the target point, estimated using a $50\,\text{cm}$ neighborhood. Finally, we include the mean intersection-over-union ($IoU$) between occupancy grids of the original and reconstructed point clouds, defined as: $IoU=|G_{\text{src}}\cap G_{\text{trg}}|\,/\,|G_{\text{src}}\cup G_{\text{trg}}|$. As in [13], the occupancy grids are computed with a voxel resolution of $20\times 20\times 10\,\text{cm}^{3}$.
@@ IV-B Compression Results and Qualitative Analysis
Figure 4: Compression results per codec. The left column presents results on SemanticKITTI (in-dataset evaluation), while the right column shows cross-dataset generalization on nuScenes, where models were trained on SemanticKITTI and applied directly to nuSces without fine-tuning. Dashed lines indicate methods that do not encode semantic labels.
@@ IV-D Ablation Studies
We further studied the effect of using predicted instead of ground-truth labels for the proposed pipeline, simulating a realistic deployment. Specifically, we tested RangeNet++ [29], KP-FCNN [14], and RandLA-Net [23] as semantic segmentation front-ends. The results, reported in Table II, show that segmentation accuracy ($mIoU$) is not directly proportional to reconstruction quality. Although RangeNet++ achieves the lowest $mIoU$, it yields the best reconstruction among the three networks, since many of its characteristic “shadow” artifacts are filtered out during clustering. In contrast, RandLA-Net, despite comparable $mIoU$, produces larger reconstruction errors due to cluster-level misclassifications, while KP-FCNN performs in between with more localized errors. These findings show that the type and distribution of segmentation errors affect compression performance, offering practical insights for deploying segmentation-compression pipelines in real systems.
@@ IV-E Downstream Robotic Tasks
Pose Graph Optimization (PGO): In a multi-agent setup, we simulate point cloud transmission between agents operating on SemanticKITTI sequences 00 and 07, followed by PGO. Each agent estimates an initial trajectory using KISS-ICP [28] Loop closure candidates between decompressed and raw scans undergo geometric verification via GICP [30]; those with registration error exceeding a predefined threshold are rejected. We evaluate two cases where pose graph edges are obtained by registering raw (00) with decompressed (07) scans and vice versa. Results are averaged over both configurations using the GTSAM [31] optimization. Performance is measured by the percent improvement in Absolute Trajectory Error (ATE) relative to ground truth. We also quantify the bandwidth required for transmitting raw vs. compressed scans at 10 Hz. As shown in Table III, MPEG and Draco suffer from high loop rejection and degraded ATE due to poor geometric fidelity, while our method closely follows raw-data trajectory accuracy with roughly 98% bandwidth reduction.
@@
Figure 6: Merged maps and zoomed views. (a) Result of KISS-Matcher [32] on KITTI [27] with the decompressed scans from sequence $0$ (using our method) and the raw scans from sequence 07. (b)–(c) Example of the decompressed frames before and after filtering, where dynamic objects (pedestrian, car, motorcyclist) are succesfully removed. (d)–(e) Side-by-side comparison of the overlapping regions: decompressed vs. raw, highlighting preserved structure.
Multi-Robot Map merging: We further evaluate the utility of our compressed representations in multi-robot map merging. Full-session maps are built from SemanticKITTI sequences using: raw-to-raw (00/07), raw 00 with decompressed 07, and the reverse. Map alignment is performed via KISS-Matcher [32], with mean Relative Transformation Error (RTE) measured against ground truth (see Fig. 6). Using our scene graph, we apply semantic pruning to remove dynamic obstacles (e.g., pedestrians or vehicles linked to road terrain) from its layered structure, improving clarity as illustrated in Fig. 6. MPEG and Draco fail to produce reliable correspondences, making them unsuitable for this task, as shown in Table III. In contrast, our method closely follows the alignment quality of raw data while reducing bandwidth by approximately 98%. These results demonstrate that our semantic scene graph-based compression preserves essential geometric and semantic fidelity, while also enabling a practical communication strategy: agents can initially exchange compressed representations, sufficient for tasks such as multi-agent pose graph optimization and map merging, and later, if needed, update with full-resolution scans when higher bandwidth becomes available.

### v2

@@ IV Experimental Evaluation
We evaluate our method’s capability to efficiently and accurately compress a point cloud scan. We compare against the commonly used codecs: Draco [draco] and the octree-based compression algorithm from Mekuria et al. [mekuria2017mpeg], which we refer to as “MPEG”. Additionally we include the current state-of-the-art approaches RENO [you2025reno] and OctAttention [octattention], as well as RecNet [stathoulopoulos2024recnet], that in contrast to ours, encodes each scan into a single latent vector through a range-image embedding. We note that the last three methods, unlike the proposed approach, or Draco/MPEG, do not encode the semantic classes.
@@ IV-A Experimental Setup
We benchmark our method on the SemanticKITTI [semkitti] dataset. In order to create the semantic scene graph we use the ground truth semantic labels and in addition we crop each point cloud to 50 meters for all methods, following common practices in the field [vizzo2023ral]. Similar to SegMap [dube2020segmap] we use sequences 05 and 06 for training and the rest (00 to 04 and 07 to 10) for the evaluation and comparison to the baselines.
Evaluation Metrics: The quality of a compression algorithm typically involves a trade-off between compression ratio and reconstruction error. To quantify compression, we use the average bits-per-point (bpp) required to store the encoded point cloud. For our method, this includes all latent vectors as well as the semantic scene graph. Following the evaluation protocol in [wiesmann2021deep], we assess reconstruction error using three metrics. The first is the symmetric point distance $D_{\text{CD}}$, defined earlier in Eq. (6) and Eq. (7), which measures the average nearest-neighbor distance between the reconstructed and ground truth point clouds. However, for many robotic tasks that rely on point-to-plane ICP registration, reconstructing the exact same points is less critical than ensuring that points lie on the correct surfaces. To address this, we also report the symmetric plane distance $D_{\perp}$, which is computed similarly to $D_{\text{CD}}$ but replaces the Euclidean distance $d(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ with the point-to-plane distance ${d}_{n}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})=|\boldsymbol{n}^{\mathsf{T}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})|$, where $\boldsymbol{n}\in\mathbb{R}^{3}$ is the ground truth surface normal at the target point, estimated using a $50\,\text{cm}$ neighborhood. Finally, we include the mean intersection-over-union ($IoU$) between occupancy grids of the original and reconstructed point clouds, defined as: $IoU=|G_{\text{src}}\cap G_{\text{trg}}|\,/\,|G_{\text{src}}\cup G_{\text{trg}}|$. As in [wiesmann2021deep], the occupancy grids are computed with a voxel resolution of $20\times 20\times 10\,\text{cm}^{3}$.
@@ IV-B Compression Results and Qualitative Analysis
Figure 4: Compression results per codec. The left column presents results on SemanticKITTI (in-dataset evaluation), while the right column shows cross-dataset generalization on nuScenes, where models were trained on SemanticKITTI and applied directly to nuScenes without fine-tuning. Dashed lines indicate methods that do not encode semantic labels.
@@ IV-D Ablation Studies
We further studied the effect of using predicted instead of ground-truth labels for the proposed pipeline, simulating a realistic deployment. Specifically, we tested RangeNet++ [rangenet], KP-FCNN [thomas2019KPConv], and RandLA-Net [hu2020randla] as semantic segmentation front-ends. The results, reported in Table II, show that segmentation accuracy ($mIoU$) is not directly proportional to reconstruction quality. Although RangeNet++ achieves the lowest $mIoU$, it yields the best reconstruction among the three networks, since many of its characteristic “shadow” artifacts are filtered out during clustering. In contrast, RandLA-Net, despite comparable $mIoU$, produces larger reconstruction errors due to cluster-level misclassifications, while KP-FCNN performs in between with more localized errors. These findings show that the type and distribution of segmentation errors affect compression performance, offering practical insights for deploying segmentation-compression pipelines in real systems.
@@ IV-E Downstream Robotic Tasks
Pose Graph Optimization (PGO): In a multi-agent setup, we simulate point cloud transmission between agents operating on SemanticKITTI sequences 00 and 07, followed by PGO. Each agent estimates an initial trajectory using KISS-ICP [vizzo2023ral] Loop closure candidates between decompressed and raw scans undergo geometric verification via GICP [fast_gicp]; those with registration error exceeding a predefined threshold are rejected. We evaluate two cases where pose graph edges are obtained by registering raw (00) with decompressed (07) scans and vice versa. Results are averaged over both configurations using the GTSAM [gtsam] optimization. Performance is measured by the percent improvement in Absolute Trajectory Error (ATE) relative to ground truth. We also quantify the bandwidth required for transmitting raw vs. compressed scans at 10 Hz. As shown in Table III, MPEG and Draco suffer from high loop rejection and degraded ATE due to poor geometric fidelity, while our method closely follows raw-data trajectory accuracy with roughly 98% bandwidth reduction.
@@
Figure 6: Merged maps and zoomed views. (a) Result of KISS-Matcher [lim2025icra-KISSMatcher] on KITTI [semkitti] with the decompressed scans from sequence $0$ (using our method) and the raw scans from sequence 07. (b)–(c) Example of the decompressed frames before and after filtering, where dynamic objects (pedestrian, car, motorcyclist) are succesfully removed. (d)–(e) Side-by-side comparison of the overlapping regions: decompressed vs. raw, highlighting preserved structure.
Multi-Robot Map merging: We further evaluate the utility of our compressed representations in multi-robot map merging. Full-session maps are built from SemanticKITTI sequences using: raw-to-raw (00/07), raw 00 with decompressed 07, and the reverse. Map alignment is performed via KISS-Matcher [lim2025icra-KISSMatcher], with mean Relative Transformation Error (RTE) measured against ground truth (see Fig. 6). Using our scene graph, we apply semantic pruning to remove dynamic obstacles (e.g., pedestrians or vehicles linked to road terrain) from its layered structure, improving clarity as illustrated in Fig. 6. MPEG and Draco fail to produce reliable correspondences, making them unsuitable for this task, as shown in Table III. In contrast, our method closely follows the alignment quality of raw data while reducing bandwidth by approximately 98%. These results demonstrate that our semantic scene graph-based compression preserves essential geometric and semantic fidelity, while also enabling a practical communication strategy: agents can initially exchange compressed representations, sufficient for tasks such as multi-agent pose graph optimization and map merging, and later, if needed, update with full-resolution scans when higher bandwidth becomes available.